Air temperature drops approximately 5.5ºF for each 1,000-foot rise in altitude above Earth’s surface (up to 30,000 ft).
a. Write a formula that relates temperature t in degrees Fahrenheit at altitude h (in thousands of feet) and a ground temperature of 65ºF. State any restrictions on h.
b. Find the temperature at 11,000 ft above Earth’s surface.

Answer by chessace(471)

t=65+5.5h, h<30.
t(11)=65+5.5*11=125.5